OpenStudy (anonymous):

given -4 is a root, find the remaining real roots for f(x)=x^3 + 2x^2 - 11x - 12

OpenStudy (anonymous):

that means \[x^3+2x^2-11x-12=(x+4)(something)\] the "something" will be a quadratic so you can find the roots of it

OpenStudy (anonymous):

you can find the something by long division ( a pain) synthetic division (very snappy) or just thinking about what it has to be

OpenStudy (anonymous):

to basically do (x+4)(something that when multiplied will equal the equation)?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

do you know how to do synthetic division?

OpenStudy (anonymous):

I don't remember much of it...

OpenStudy (anonymous):

well then lets just think synthetic division is real real easy but almost impossible two write here

OpenStudy (anonymous):

isnt it where you write down all the coefficients and then multiply the numbers by the root?

OpenStudy (anonymous):

\[x^3 + 2x^2 - 11x - 12=(x+4)(ax^2+bx+c)\] it should be pretty clear that it will be \[x^3+2x^2-11x+12=(x+4)(x^2+bx-3)\] right? we just don't know \(b\) yet

OpenStudy (anonymous):

yes it is multiply and add etc

OpenStudy (anonymous):

but if you got how i wrote the second line above, we can find \(b\) real easy

OpenStudy (anonymous):

when i did synthetic division, i got what seems to be a remainder of 40...

OpenStudy (anonymous):

Im not sure exactly how to do it the other waay...

OpenStudy (anonymous):

then you screwed up because the remainder has to be zero lets find \(b\) did you understand how i got \[x^3+2x^2-11x+12=(x+4)(x^2+bx-3)\]

OpenStudy (anonymous):

 sorry for the bad writing but this is what i did..